OpenStudy (anonymous):

How to find the arc length of y=1+6x^(3/2), 0

OpenStudy (amistre64):

measure the hypotenuse of a very tiny triangle from 0 to 1

OpenStudy (amistre64):

\[\int_{D}^{}ds=\int_{0}^{1}\sqrt{1+[f'(x)]^2 }dx\]

OpenStudy (amistre64):

y=1+6x^(3/2) y' =9x^(1/2) i think

OpenStudy (amistre64):

\[\int_{0}^{1}\sqrt{1+(9x^{1/2})^2 }dx\] \[\int_{0}^{1}\sqrt{1+81x}\ dx\]

OpenStudy (anonymous):

Thank you, I appreciate your help!

OpenStudy (amistre64):

youre welocme, with any luck, that made sense :)

OpenStudy (amistre64):

and was typed up correctly lol